Commander Report – March 31st

Crew Commander Report by Otsmar Villarroel/Crew 177 Commander
March 31st/2017
Today, March 31st is our mission’s last day.  This week has been definitely unique.  Thanks to McLennan Community College, Mars Desert Research Station, and my crewmates for this opportunity.
Today afternoon was used to conduct an EVA led by Pitchaya Jingjit.  She collected twenty-one different soil samples from five different places of interest around the MDRS location with the purpose of bringing back to McLennan Community College, in Waco-Texas in order to find bacteria producing antibiotics to combat antibiotic resistance crisis.  Jingjit was accompanied by Caleb Li, Esteban Ramirez, and Otsmar Villarroel.
Geologists, Espinoza and Li, worked in the science lab during the afternoon.
They use the regolith samples collected during geology EVAs and performed Sieving analysis on each of their samples.  They use different trays, each with a different mesh size.  Back on McLennan Community College campus, they are planning to perform spot test analysis for the following: Carbonate properties, oxide, iron, Uranium, and Potassium.
Dinner consisted of salmon and mashed potatoes.  After dinner crew members started cleaning the MDRS facilities and already started to pack and getting ready to return to Earth with the satisfaction of being part of this experience.
See you later Mars! And one more time, thanks to everybody that makes it an awesome experience.

Commander Report – March 30th

Sol 5/March 30th, 2017
Crew Commander Report by Otsmar Villarroel, Crew 177 Commander
Although today was not as beautiful as previous days with a beautiful blue sky and calm wind, it was a great day for the crew’s mission here on the red planet. After breakfast, a group consisting of Parker, Li, Espinoza, and Jingjit left the Hab to conduct some geology experiments.
The main goal of this EVA was to study the terrain from a geological point of view and thus collect some geology data of interest.  The location for this EVA was chosen in order to study different kinds of geologic structures and the methods by which they were created. The pattern of lithification in the sedimentary rock field was unique and interesting to the crew.
Former evidence of a vigorous fluvial environment was observed at the second location. Determination of width and height were performed during the EVA. After collection of the data, the Geology EVA group returned to the Hab.
After enjoying lunch, at about 2:30 pm, Martian time, Crew members LaBarre and Espinoza, had the opportunity of testing her Emergency Medical Service Rover (EMRS) outside the Hab. The EMSR is a robot designed and programmed to retrieve an injured astronaut out on an EVA at Mars and bring him or her back to the Hab for medical assistance. During the test LaBarre and Espinoza detected some operational/mechanical issues.  However, they plans correcting them and a new test, pending approval, will be performed tomorrow.
Engineer, Espinoza, completed their daily routine without anomalies to be reported.  Greenhab officer, Ramirez, worked on taking care of plants as well as on his individual project, the bike generator.  He is planning to collect final results tomorrow to be analyzed back in Waco, Texas.  Our microbiologist, Jingjit, began and finished the gram stain process for the bacterial samples taken previously and attempted to use the microscope. Crew member, Li, spent part of the afternoon working on his clock.
After dinner, crew spent the crew spent a few hours helping coming up with a marketing plan to promote MDRS at our community college, McLennan Community College.  Tomorrow, pending approval, an EVA is planned. The purpose of this EVA is the collection of soil samples surrounding the Mars Desert Research Station (MDRS) to bring back to McLennan Community College in Waco, Texas in order to find bacteria producing antibiotics to combat antibiotic resistance crisis.
Last night, our crew astronomer, Becky Parker, used the telescope, with the main goal of completing a two-star alignment as well as a four- star calibration procedure. Unfortunately, while she was able to have the telescope working properly, she could not get the telescope to focus as expected.  We are planning to contact Mission Support looking for recommendations.

Commander Report – March 29th

Commander Report
Sol 5
March 29,2017
Otsmar Villarroel, Crew 177 Commander
Although today’s temperature was slightly higher than previous days, the day here on Mars has been beautiful with a spectacular blue sky.
All six crewmembers were up by 0700 local time as requested by the crew Commander the night before.  The crew enjoyed together scrambled eggs with sausage and biscuits for breakfast.  A few minutes after breakfast, a group consisting of Espinoza, Espinoza, Li, and Parker was sent out to Candor Chasma for a second attempt of looking for the presence of water.
 While during this effort, as the first one few days ago, the presence of water was not detected, the crew found the presence of moisture in some areas by Candor Chasma.  It was noticed after digging into the sand in the bottom of the chasma. Also, the presence of moisture was observed underneath some rocks.  Results are promising and further research regarding this topic has to be done.
After 2 hours 30 minutes, approximately, by noon (Martian time) the expediters returned to the Hab.  All eight crewmembers enjoyed lunch together.  Lunch was prepared crew members Quaas and Jingjit.
Part of the afternoon was used by the crew members to work on their independent projects.  Crew members have found some difficulties with their projects and they are making plans to solve them as soon as possible before rotation is over.  Also, crew member Jingjit continued working on her microbiology experiments.
As usual, dinner time was used to spend time together and discuss tomorrow’s plans.  Today it is crew member LaBarre’s birthday and we celebrated it with a birthday cake. The crew is planning to complete the geology EVA, pending approval. This would be the second half of EVA #4 that was shortened due to NPR reporter.

Commander Report – March 28th

Four sols have passed since The McLennan Community College Highlanders Crew 177 arrived to Mars.  For sure, four sols full of challenges and amazing experiences.  Every activity around here is designed with the main purpose of providing us a unique experience other than that offered by a traditional classroom environment.

After all eight crewmembers took breakfast together, at about 0845 local time; Crew 177 got visited by Rae Bichell from NPR.  Rae Bichell spent most of the day with us learning about our routine here on Mars. She is was very interested in knowing about our experiences, projects, our college as well as all related with the MDRS program.  After giving her a short talk about our goals followed by a tour around the MDRS complex, she was invited to escort us during our morning Geology EVA.

After the daily routine of getting suited followed by the depressurization  process before going out of the Hab, the Geology EVA crew departed at about 1045 local time.  The EVA group consisted of Elijah Espinoza, Pitchayapa Jingjit, Esteban Espinoza, and Becky Parker. The geology experiments conducted today consisted of assessing many different regolith characteristics in the area around the Hab. Understanding what this material is made out of helps to determine what resources could be available. Also, crew members learned about the use of a Brunton compass to measure as strike and dip, both useful geological parameters which provide valuable information about what angle a particular geological formation has been uplifted.

While EVA team were collecting samples and taking geological measurement outside the Hab, remaining crew stayed in the Hab serving as communication support.  During that time they also prepared lunch which was enjoyed by the crew members and guest.  Menu included delicious teriyaki chicken, rice, corn, and red pepper.

During the afternoon, Pitchayapa started working on her microbiology experiments.  The point of her experiments is to swab the Hab for the presence of Gram negative and enteric bacteria and Gram positive staph bacteria.  Both have health implications.

Engineer completed their daily checking routine without major anomalies.

During dinner time the crew discussed the plan for tomorrow.  While some crew members showed some signs of fatigue during the day, Crew 177 members are looking forward to keep doing their best in order to successfully complete this mission.  The challenges make us stronger!
Tomorrow, pending approval, will consist of a morning chemistry EVA with Elijah, Esteban, Caleb, and Becky.  Rest of the afternoon will be spend working on individual projects and microbiology experiments conducted by Pitchayapa.

Good night from Mars

Commander REport – March 27th

Commander Report Crew 177:
Otsmar Villarroel, Commander
Sol 2
Earth date: 3/27/2017
Today the crew awoke around 7:00 am and enjoyed the morning playlist courtesy of Caleb. Lead Engineer, Elijah Espinoza, took care of filling the water tanks for our future water consumption in the HAB.  We are very glad that water issue was solved.

After breakfast, the morning EVA left the hab at 9:11 am for Candor Chasma to complete the first Chemistry EVA.  Commander Otsmar, Engineer Elijah, Chemist Joseph, Biologist Pitchayapa, and I spent most of the morning in the canyon trying to find water.  On Earth, this molecule, made up of three atoms, is many times associated with life. While expeditioners could not find the presence of this vital molecule in its more stable state, under normal conditions, it was evident that water was part of that beautiful scene some time ago.  The existence of beautiful geological structures, presumably sculpted by the power of water, were observed along with the presence of gypsum, CaSO4*2H2O, which support our ideas.  Some unidentified animals’ (we hope) footprints were observed as well as some plants, trees, flowers, and cacti. We also discovered an interesting cave that showed signs of nesting by some unknown creature.  Crew returned safely at approximately 11:31 am.  Right after arriving, Elijah checked the ATVs gas level and then proceeds to fill them.

After taking lunch together, the crew spent the afternoon doing individual research and reviewing our media training in preparation for the NPR reporter that will be spending part of the day with us tomorrow.  We all are very excited about that visit and ready to share our experiences here at MDRS as well as talk about our mission and projects.

Engineer, Elijah, went out to complete his daily routine, making sure everything is working as expected.

After dinner, crew discussed about tomorrow’s plan.

Otsmar Villarroel
177 Crew Commander

Commander Report – March 26th

Crew 177 Commander’s Report 26 March 2017 Otsmar Villarroel Commander Report

26 March 2017 Sol 1

Today we officially entered sim at approximately 10:40 am after saying goodbye to the Poland Crew 176. After getting the space suits on and following the pressurization protocols, at about 1115 local time first EVA group left the HAB to complete their day’s EVA 1 duties.  We conducted two orientation EVAs. First group was led by Otsmar Villarroel while second group by Becky Parker. Both EVA groups (morning and afternoon) gained experience driving ATVs as well as using GPS navigation while staying in sim. Also both EVAs focus on helping orient crew members to their surroundings, learn major landmarks along the main ATV road Besides discovering the presence of another form of life other than us on this planet- Martian cows- crew members observed:

  • Ancient sand dunes
    • Caprock erosion from wind and water (mass wasting)
    • A mushroom shaped hoodoo
    • A creek bed exhibiting signs of efflorescence
    • A field of gypsum bordering a substantial quantity of bivalve fossils

During EVA 1, the crew remaining in the hab cleaned and organized the hab, completed and updated the food inventory and took care of lunch.  Also, both Elijah and Victoria started assembling their rover after taking it apart for travel.  The crew came together for lunch.


Second group left the HAB at 1510 local time.  During EVA 2, the crew remaining inside the hab worked on reports, research, updated the first aid kit inventory and made some bread.  Esteban spent part of the afternoon re-assembling his power generator after taking it apart for travel.  After arriving from second EVA, Engineer Elijah Espinoza completed his routine engineering round.


Dinner menu includes macaroni and cheese, mashed potatoes, and broccoli prepared by Head Chef Quaas with the assistance of Victoria LaBarre. Our plan for tomorrow includes media training, independent research, and, pending approval, a Chemistry EVA in the morning to Candor Chasma.

Crew members feel motivated and ready to face more challenges here on Mars.  However, the lack of water in the HAB has become a big issue keeping the whole crew worry.  Crew members requested more water and we expect this issue get solved anytime soon.  We are doing our best to save as much water as we can, but it is not an easy task for a crew of 8 people.
Crew 177
Crew Commander
Otsmar Villarroel

Commander Report – March 25th

Crew 176 Commander’s Report 25 March 2017
Natalia Zalewska
Commander Report
25 March 2017 – Sol 14

Dear Mission Support,

This was our last day at MDRS. We have EVA scheduled for this afternoon,
however, due to the poor weather conditions and the workload related to
packing up, we decided to cancel it and finish the simulation today.

In the morning, we worked on the video recordings. Green Hab Officer the
two crew members went to Grand Junction to pick up a rental car we will
use to transport the robot and other equipment. We finished cleaning the
upper and lower deck as well as helmets. In the afternoon we hosted Crew
177. We were happy to explain our guests how to run and maintain the MDRS
campus as well as what are the rules to follow when using ATVs and
engaging in EVAs. The evening was particularly long as we needed to finish

The crew is in a great mood, as it always is. We are also grateful to Mars
Society for the Spirit and Opportunity we have been provided at MDRS!

Best regards,

Natalia Zalewska
Commander, MDRS Crew 176



Commander Report Crew 177:
Otsmar Villarroel
Sol 1
Earth date: 03/25/2017

Good evening!

Today we met the very helpful Crew 176 at MDRS.  Upon arrival at the hab, crew members were trained by Crew 176 members.  Training included a brief tour around the MDRS facilities as well as over use of the ATVs, operation report data, water leak issue, etc.  After the training we went to Bull Mountain Market to pick up our food according to the inventory sent by crew 176 as requested by Shannon.

We’ve made a plan for tomorrow, which will include, pending approval, our Orientation EVAs. We are expecting to enter sim tomorrow morning at 9:30 am, if possible.

From Crew 177; Hanksville, UT
Crew Commander

Commander’s Report – March 23rd

Crew 176 Commander’s Report 23 March 2017
Natalia Zalewska
Commander Report
23 March 2017 – Sol 12

Dear Mission Support,

While we are heading to the end of our mission, we continue to maintain
our daily routine. Thus, we woke up today in the morning, had a breakfast
together as well as some of us completed psychological tests. We made sure
holter monitors work but due to the increasing fatigue of the entire team,
we skipped the sport activities.

Despite poor weather conditions, we undertook EVA as planned. This time
the crew members stayed in the close proximity of the habitat without the
need to use ATVs. The main focus was on the use of the rover Ares 2,
including in the role of the support tool when collecting geological

The Crew Engineer spent a significant amount of time working on the rover
and a corresponding trailer, and the Green Hab Officer was happy to
discover new plants in the pots. Also, just as every day, the entire crew
worked on preparing audio-visual data for both education and outreach

Starting from tomorrow, we will work on packing our staff, cleaning the
habitat and we will get mentally ready to leave. Our mood and teamwork is
still very good.

Best regards,

Natalia Zalewska
Commander, MDRS Crew 176

Commander Report – March 21st

Crew 176 Commander’s Report 21 March 2017
Natalia Zalewska
Commander Report
21 March 2017 – Sol 10

Dear Mission Support,

Our daily routine remained unchanged today. We woke up in the morning and
had a breakfast together, oats with dried fruits, as usual. Afterwards
some of us completed psychological tests and then the selected members of
the crew got ready for EVA and left. EVAs continue to be the most
interesting tasks we undertake at MDRS in terms of research, technology
and documentation-related activities and constitute one of the main
subjects of our discussions. Personally, I find EVAs a rare opportunity
for undertaking very interesting geological research.

We had lunch together and took some rest. In the afternoon we proceeded
with working on the audio-visual content for the mass media, including
taking photographs and video recordings in the MDRS hab. We were glad to
have the Internet connection back (at least for a while) and online
communication with Mission Support re-established. Also, the Green Hab
Officer was happy to learn there are two more plants visible in the pots,
all of the belonging to the Zinnia flowers family.

We expect to have the Internet issues fully solved by tomorrow. We will
proceed with EVA and documentation activities, as well as my analysis of
geological samples. We will also continue to test the robot. We are aware
that the weather conditions will deteriorate but we hope it won’t
affect our EVAs.

Bets regards,

Natalia Zalewska
Commander, MDRS Crew 176